calculate the concentration of each of the ions in a mixed solution in which no reaction occurs when 125 mL of 0.0240 M strontium sulfide combines with 75.0 mL of 0.0480 M strontium chloride.
1
Expert's answer
2020-03-04T07:59:22-0500
  1. Find moles of SrS
"0.125 L\times\frac{0.0240 mol}{L}=0.003 mol"



"SrS(aq) \rightarrow Sr^{2+}(aq)+S^{2-}(aq)"


"n(SrS)= n(Sr^{2+})=n(S^{2-})=0.003 mol"


2.Find moles of SrCl2


"0.075 L \times \frac{0.0480 mol}{L} = 0.0036 mol"



"SrCl_2(aq)\rightarrow Sr^{2+}(aq) + 2Cl^-(aq)"



"n(SrCl_2)=n(Sr^{2+})=0.0036 mol"


"n(Cl^-)=0.0036 mol(SrCl_2)\times \frac{2 mol(Cl^-)}{1mole (SrCl_2)}=0.0072 mol(Cl^-)"

3.Find total volume of a new solution:


"V_{total} = 0.125L +0.075 L = 0.2 L"


4.Find concentration of each of the ions


"c(S^{2-})=\frac{0.003 mol}{0.2 L}= 0.015 \frac{mol}{L}"


"c(Cl^-)=\frac{0.0072 mol}{0.2 L}= 0.036 \frac{mol}{L}"


"c(Sr^{2+})=\frac{0.003mol+0.0036mol}{0.2 L}= 0.033 \frac{mol}{L}"